QUESTION FIVE


From the figures given below;


OUTPUT(UNITS) TOTAL REVENUE (K’000)


50 750


60 840


70 910


80 960


90 990


100 1000


You are required to:


a) Calculate the average revenue [3 marks]


b) Calculate the marginal revenue [2 marks]


c) Draw the average revenue curve [2 marks]


d) Determine the market structure [1 mark]


e) Mention two features of the market structure in (d) [2 marks]

1
Expert's answer
2022-02-09T10:53:00-0500

a) Calculate the average revenue [3 marks]

The formula to calculate average revenue is given by:

Average revenue = "\frac{Total\,revenue}{Total\,units}"

For 50 units

Average revenue = "\frac{750}{50} = 15"

For 60 units

Average revenue = "\frac{840}{60} = 14"

For 70 units

Average revenue ="\frac{910}{70} = 13"

For 80 units

Average revenue ="\frac{960}{80} = 12"

For 90 units

Average revenue ="\frac{990}{90} = 11"

For 100 units

Average revenue ="\frac{1000}{100} = 10"


The pictorial representation is shown below:



b) Calculate the marginal revenue [2 marks]

Marginal Revenue = Change in Revenue / Change in Quantity


For 60 units

Marginal revenue = "\frac{840-750}{60-50} = 9"

For 70 units

Marginal revenue ="\frac{910-840}{70-60} = 7"

For 80 units

Marginal revenue ="\frac{960-910}{80-70} = 5"

For 90 units

Marginal revenue ="\frac{990-960}{90-80} = 3"

For 100 units

Average revenue ="\frac{1000-990}{100-90} = 1"



c) Draw the average revenue curve [2 marks]



d) Determine the market structure [1 mark]

The name of this market structure is Monopoly. Because the marginal revenue curve in this situation is downward sloping.

e) Mention two features of the market structure in (d) [2 marks]

  • Barriers to entry into the market exist.
  • In the market, there is just one seller.
  • Discrimination based on price